The demand for ice cream cones is P=1600 AND Qd is 2
The supply of ice cream cones is P=400 and Qs is 1. The price of a cone is expected in cents , and the quantities are expressed in cones per day. To find the Equilibrium price (P*) and the equilibrium quantity (Q*), substitute Q* for QD and QS and P* for P.
(For QD, P=a-bQd and for QS, P=c+dQs)
1
Expert's answer
2019-03-18T11:19:45-0400

Pd = 1600, Qd = 2

Ps = 400, Qs = 1.

For QD, P=a - bQd and for QS, P=c+dQs, so:

1600 = a - 2b, 400 = c + d,

In equilibrium a - bQ = c + dQ,

Q = (a - c)/(b + d).